XxEmilyxX:

Line MN passes through points M(4, 3) and N(7, 12). If the equation of the line is written in slope-intercept form, y = mx + b, what is the value of b? –15 –9 3 9

supie:

Ok so what we have to do is find the slope. \[(12 - 3) / (7 - 4) = 9 / 3 = 3\] \[3=3\ is\ true\ so\ m=3\] and \[y = 3x + b\] but when \[y = 3\]\[x = 4\] So we plug in the numbers from the variables \[3 = 3 (4) + b\]\[b + 12 = 3\] \[\ b\ =\ ?\]